# Exact Random Graph Matching with Multiple Graphs

### v1

@@ 3 Main Results and Algorithm
There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies
@@
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1).$
Theorems 2 and 3 together characterize the threshold for exact recovery. A few remarks are in order.
  1. 1.
    For $m=2$, the condition $Cs(1-(1-s)^{m-1})>1$ reduces to $Cs^{2}>1$, which is known to be necessary and sufficient for exactly matching two graphs [CK17, WXY22].
  2. 2.
    For any $m>2$, there is a non-empty region in the parameter space defined by
    $Cs(1-(1-s)^{m-1})>1>Cs^{2}.$
    For any $C$ and $s$ in this region, it is impossible to exactly match any two graphs $G_{i}$ and $G_{j}$ without using the other $m-2$ graphs as side information. Upon using them, however, it is possible to exactly match all nodes across the $m$ graphs. This is illustrated in Figure 2.
@@
Figure 2: Regions in parameter space. Orange: Exactly matching $m$ graphs is impossible even with $m$ graphs. Blue: Exactly matching $2$ graphs is possible with $2$ graphs. Striped: Impossible to match $2$ graphs using only the $2$ graphs, but possible using $m$ graphs as side information.
3.1 Algorithms for exact recovery
For any two graphs $H_{1}$ and $H_{2}$ on the same vertex set $V$, denote by $H_{1}\vee H_{2}$ their union graph and by $H_{1}\wedge H_{2}$ their intersection graph. An edge $\{i,j\}$ is present in $H_{1}\vee H_{2}$ if it is present in either $H_{1}$ or $H_{2}$. Similarly, the edge is present in $H_{1}\wedge H_{2}$ if it is present in both $H_{1}$ and $H_{2}$.
A natural starting point is to study the maximum likelihood estimator (MLE) because it is optimal. To that end, we compute the log-likelihood function; the details are deferred to Appendix A.
Theorem 4.
Let $\pi_{12},\cdots,\pi_{1m}$ denote a collection of permutations on $\{1,\cdots,n\}$. Then
$\log\mathbb{P}\left(G_{1},\cdots,G_{m}\ |\ \pi^{*}_{12}=\pi_{12},\cdots,\pi^{*}_{1m}=\pi_{1m}\right)\propto\text{const.}-\left|E\left(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}\right)\right|,$
where const. depends only on $p,s$ and $G_{1},\cdots,G_{m}$.
Theorem 4 reveals that the MLE for exactly matching $m$ graphs has a neat interpretation: simply pick $\pi_{12},\cdots,\pi_{1m}$ to minimize the number of edges in the corresponding union graph. This is presented as Algorithm 1. Despite this nice interpretation of the MLE, its analysis is quite cumbersome. We instead present and analyze a different estimator, presented as Algorithm 2.
require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$
1 for $(\pi_{12},\pi_{13},\cdots,\pi_{1m})$ such that each $\pi_{1j}$ is a permutation on $[n]$ do
2       $W(\pi_{12},\cdots,\pi_{1m})\leftarrow|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|$
3
4 end for
return $(\widehat{\pi}_{12}^{\text{ML}},\cdots,\widehat{\pi}_{1m}^{\text{ML}})\in\operatorname*{arg\,max}_{\pi_{12},\cdots,\pi_{1m}}W(\pi_{12},\cdots,\pi_{1m})$
Algorithm 1 Maximum likelihood estimator
require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$, Integer $k$
// Step 1: Pairwise matching
@@ 3.1 Algorithms for exact recovery
2       $\widehat{\nu}_{ij}\leftarrow\operatorname*{arg\,max}_{\pi}|\mathsf{core}_{k}\left(G_{i}\wedge G_{j}^{\pi}\right)|$
       $\widehat{\mu}_{ij}\leftarrow\widehat{\nu}_{ij}\text{ with domain restricted to}\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$
        // $k$-core estimator
3

### v2

@@ 3.1 Exact recovery threshold: Positive results
There exists an estimator, namely Algorithm 1 described in the next subsection, whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1).$
A few remarks are in order.
  • The sufficient condition (1) of Theorem 1 can be slightly improved to the condition
    $\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}-\max_{j\in[n]}e^{-d_{j}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$ (2)
    Let $v^{*}\in\operatorname*{arg\,min}_{i\in[n]}d_{i}$. It can be shown that Algorithm 1 applied to the graphs $G_{k}$ yields that all nodes except $v^{*}$ can be exactly matched. The node $v^{*}$ is then the only unmatched node and can be manually matched without ambiguity.
  • Rácz and Sridhar (RS, 23) studied the graph alignment problem when $m=2$. Setting $m=2$ in (1) yields $\sum_{i=1}^{n}e^{-d_{i}s^{2}}=o(n^{-\alpha})$, which is equivalent to their sufficient condition.
  • Algorithm 1 does not require access to the parameter matrix $\mathbf{P}$.
The sufficient condition (1) can be specialized to various models that are captured by inhomogeneous random graphs. Some of these are presented next.
Corollary 2 (Correlated SBMs)
Let $G_{1},\cdots,G_{m}$ be correlated SBMs obtained from the subsampling model. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $q_{ab}=o(n^{\alpha/2-\varepsilon-1})$ for all $a$ and $b$. Suppose that
$\sum_{a=1}^{M}|V_{a}|\cdot\exp\left(-s\left(1-(1-s)^{m-1}\right)\cdot\sum_{b=1}^{m}|V_{b}|q_{ab}\right)=o(n^{-\alpha}).$
@@
There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
Corollary 3 (Correlated RGGs)
Let $G_{1},\cdots,G_{m}$ be correlated RGGs obtained from the subsampling model. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p=o(n^{\alpha/2-\varepsilon-1})$. For a node $i$, let $N_{r}(i)$ denote the number of nodes in its $r$-neighborhood, i.e. $N_{r}(i)\triangleq|\{j:\|x_{i}-x_{j}\|_{2}\leq r\}|$. Suppose that
$\sum_{i=1}^{n}e^{-ps(1-(1-s)^{m-1})\cdot N_{r}(i)}=o(n^{-\alpha}).$
There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
Corollary 4 (Correlated CLGs)
Let $G_{1},\cdots,G_{m}$ be correlated CLGs obtained from the subsampling model. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $w_{i}=o(n^{\alpha/2-\varepsilon-1})\sqrt{\sum_{k\in[n]}w_{k}}$ for all $i$. Suppose that
$\sum_{i=1}^{n}e^{-w_{i}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$
There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
Next, we show that the sufficient condition (2) can be improved for a class of parameter matrices if the algorithm has access to $P$.
Theorem 5
Suppose $\varepsilon$ and $s$ are fixed with $\varepsilon>0$ and $0<s\leq 1$ If $d$ varies with $n$ such that:
$ds\geq(1+\epsilon)\log n,$ (3)
then there exists a parameter matrix $P$ with $d_{i}\triangleq\sum_{j\neq i}P_{ij}\leq d$ for $i\in[n]$ and $p_{\max}\triangleq\max_{i,j}P_{i,j}=\frac{\log^{2}(n)}{n}$ such that exact recovery is possible, i.e. $m$ graphs $G_{1},G_{2},\cdots,G_{m}$ obtained by subsampling each edge independently with probability $s$ from a parent graph $G\sim\mathrm{RG}(n,P)$ can be exactly matched with probability $1-o(1)$ by some algorithm that has access to $P$.
Taking $ds=(1+\epsilon)\log n$ in Theorem 5 yields the following corollary.
Corollary 6
There exists a parameter matrix $P$ satisfying
$\sum_{i=1}^{n}e^{-d_{i}s(1-2\varepsilon)}-\max_{j\in[n]}e^{-d_{j}s(1-2\varepsilon)}=\Omega(n^{\varepsilon}),$ (4)
such that exact recovery is possible by some algorithm that has access to $P$.
Remark 7
Corollary 6 shows that if $(1-s)^{m-1}>2\varepsilon$, then the sufficient condition (2) of Theorem 1 does not hold for some $P$, but exact recovery is nevertheless possible. Roughly speaking, the matrix $P$ itself is similar to an additional graph that can help in matching by some algorithm that has access to $P$.
3.2 Exact recovery threshold: negative results
While Corollary 6 shows that (2) is not a sharp sufficient condition for general parameter matrices, we prove that this condition is sharp in the homogeneous case of Erdős-Rényi graphs. This provides the sharp threshold for exactly matching $m$ ER graphs.
Theorem 8
Let $C>0$ and $p_{ij}=C\log n/n$ for each $i<j$, and let $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose that
$\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}=\Omega(1),$ (5)
or equivalently that $Cs(1-(1-s)^{m-1})<1$. Then exact recovery is impossible, i.e. the output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ of any estimator satisfies
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-\Omega(1).$
In the setting of ER graphs with $p_{ij}=C\log(n)/n$, Theorems 1 and 8 establish that exact recovery is possible if and only if $Cs(1-(1-s)^{m-1})>1$. When $m=2$, this condition reduces to $Cs^{2}>1$, recovering the reconstruction threshold of Cullina and Kiyavash (CK, 17). Thus, when $m>2$, there is a regime given by
$Cs(1-(1-s)^{m-1})>1>Cs^{2},$
where exact recovery is impossible using only the two graphs but possible using all $m$ graphs. This is illustrated in Figure 2.
@@ 3.2 Exact recovery threshold: negative results
Figure 2: Regions in parameter space for the homogeneous setting. Orange: Exactly matching $m$ graphs is impossible even with $m$ graphs. Blue: Exactly matching $2$ graphs is possible with $2$ graphs. Striped: Impossible to match $2$ graphs using only the $2$ graphs, but possible using $m$ graphs as side information.
We turn next to negative results in the heterogeneous case. To that end, consider the following lemma about isolated nodes in an inhomogeneous random graph.
Lemma 9
Let $H\sim\mathrm{RG}(n,\mathbf{Q})$, and let $r_{i}=\sum_{j\neq i}q_{ij}$. Suppose that $q_{\max}=o(1/\sqrt{n})$ and
$\left(\sum_{i=1}^{n}e^{-r_{i}}\right)-\max_{j\in[n]}e^{-r_{j}}=\Omega(1).$ (6)
With probability bounded away from 00, there are at least two isolated nodes in $H$.
Lemma 9 is proved in Appendix B.1. Applying it to $H=G_{1}$ yields a regime where there are at least two isolated nodes $u$ and $v$ in $G_{1}$, with probability bounded away from 00. Since it is impossible to distinguish between $u$ and $v$ in $G_{1}$, it is impossible to match them with probability greater than $1/2$. Exactly matching the $m$ graphs is therefore impossible in this regime. This implies the following converse to Corollary 6.
Theorem 10
Consider the model of $m$ correlated graphs with parameters $P$ and $s$. Let $d_{i}=\sum_{j\neq i}p_{ij}$. If $p_{\max}=o(1/\sqrt{n})$ and
$\left(\sum_{i=1}^{n}e^{-d_{i}s}\right)-\max_{j\in[n]}e^{-d_{j}s}=\Omega(1),$ (7)
then for any algorithm the probability of exactly matching all $m$ graphs is $1-\Omega(1)$.
Remark 11
If we ask for a given $P,s,m$ whether exact recovery is possible by some algorithm, then implicitly algorithms with access to $P$ should be included because $P$ is deterministic. It remains an open problem to give the sharp recovery threshold for a given general $P,s,m$. Theorem 1 provides a performance floor while Theorem 10 provides a performance ceiling. Theorem 8 shows Theorem 1 is tight for homogeneous $P$ while Corollary 6 shows Theorem 10 is tight for some choices of $P$.
Remark 12
Lemma 9 can also be used to give a second proof of the converse result in the homogeneous setting (i.e. Theorem 8) in the special case of $m=2$. Setting $H=G_{1}\wedge G_{2}^{\prime}\sim\mathrm{RG}(n,\mathbf{P}s^{2})$ in Lemma 9, we have that if $p_{\max}=o(1/\sqrt{n})$ and
$\left(\sum_{i=1}^{n}e^{-d_{i}s^{2}}\right)-\max_{j\in[n]}e^{-d_{j}s^{2}}=\Omega(1),$ (8)
then with probability bounded away from 00, there are at least two nodes $u$ and $v$ that are isolated in $G_{1}\wedge G_{2}^{\prime}$, i.e. $u$ and $v$ both have disjoint sets of neighbors in $G_{1}$ and $G_{2}^{\prime}$. This implies that the maximum likelihood for Erdős-Rényi graphs, given in Appendix A, fails. In the heterogeneous setting, however, the existence of two isolated nodes in the intersection graph $G_{1}\wedge G_{2}^{\prime}$ does not imply impossibility of exact matching, because the maximum likelihood estimator no longer has this simple form.
3.3 Estimation Algorithm
The algorithm that achieves the guarantee of Theorem 1 is based on the idea of transitive closure: First, the graphs are matched pairwise to obtain partial matchings $\widehat{\mu}_{ij}$ between each pair of graphs. These matchings can then be boosted via transitivity. The idea is formalized in Algorithm 1, which runs in two steps: In step 1, the $k$-core estimator, for a suitable choice of $k$, is used to pairwise match all the graphs. For any $i$ and $j$, the $k$-core estimator selects a permutation $\widehat{\nu}_{ij}$ to maximize the size of the $k$-core of $G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}}$ (the $k$-core of a graph $G$ is the largest subset of vertices $\mathrm{core}_{k}(G)$ such that the induced subgraph has minimum degree at least $k$). It then outputs a matching $\widehat{\mu}_{ij}$ by restricting the domain of $\widehat{\nu}_{ij}$ to $\mathrm{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$. Note that the matchings $\widehat{\mu}_{ij}$ need not be complete matchings. In step 2, these partial matchings are boosted as follows: If a node $v$ is unmatched between two graphs $G_{i}$ and $G_{j}$, then search for a sequence of graphs $G_{i},G_{k_{1}},\cdots,G_{k_{\ell}},G_{j}$ such that $v$ is matched between any two consecutive graphs in the sequence. If such a sequence exists, then extend $\widehat{\mu}_{ij}$ to include $v$ by transitively matching it from $G_{i}$ to $G_{j}$. We emphasize the following characteristics of the algorithm.
require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $[n]$, Integer $k$
// Step 1: Pairwise matching using the $k$-core estimator
@@ 3.3 Estimation Algorithm
2       $\widehat{\nu}_{ij}\leftarrow\operatorname*{arg\,max}_{\pi}|\mathrm{core}_{k}\left(G_{i}\wedge G_{j}^{\pi}\right)|$
3       $\widehat{\mu}_{ij}\leftarrow\widehat{\nu}_{ij}\text{ with domain restricted to}\mathrm{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$
4